A number consists of two digits. The number is equal to five times the sum of its digits. If you add 9 to the number, the order of the it's digits is reversed.What is the number?
Answer by jorel1380(3719)

Let m and n be the two digits. Then:
10m+n=5(m+n) and
10m+n+9=10n+m
So:
10m+n=5m+5n
5m=4n
5m/4=n
And
10m+5m/4+9=50m/4 +m
40m+5m+36=50m +4m
9m=36
m=4
n=5
The number is 45
☺☺☺☺